40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 180 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  8037   Fri Feb 8 15:53:48 2013 JenneUpdateLockingPRMI work

Quote:

I completely agree with Koji.  We definitely should have locked the half PRC first.  We were all set up for that. 

 I reminded Jamie this morning that we were not, in fact, set up yesterday for a half PRC.  I had extracted what was the flat test mirror, to put in as PR2.  The test mirror was the better of the 2 G&Hs that we had measurements for, so I had used it as the flat test mirror, but then also wanted it to be the more permanent PR2.  After doing the PR2 flip, the IFO was naturally all aligned for PRMI, which is part of why we just did that.

Anyhow, Jamie used his tallness to put the other measured G&H mirror into the mount, and put that in front of the BS.  He aligned things such that he saw fringes in the half PRC. 

I then aligned POP onto the camera, and onto the PD.  Yuta is confirming that we're maximally on the REFL PDs.

We're starting locking in 5 min.

  8036   Fri Feb 8 12:43:26 2013 yutaUpdateComputersvideocapture.py now supports movie capturing

I updated /opt/rtcds/caltech/c1/scripts/general/videoscripts.py so that it supports movie capturing. It saves captured images (bmp) and movies (mp4) in /users/sensoray/SensorayCaptures/ directory.
I also updated /opt/rtcds/caltech/c1/scripts/pylibs/pyndslib.py because /usr/bin/lalapps_tconvert is not working and now /usr/bin/tconvert works.
However, tconvert doesn't run on ottavia, so I need Jamie to fix it.

videocapture.py -h:
Usage:
    videocapture.py [cameraname] [options]

Example usage:
    videocapture.py MC2F -s 320x240 -t off
       (Camptures image of MC2F with the size of 320x240, without timestamp on the image. MUST RUN ON PIANOSA!)
    videocapture.py AS -m 10
       (Camptures 10 sec movie of AS with the size of 720x480. MUST RUN ON PIANOSA!)


Options:
  -h, --help          show this help message and exit
  -s SIZE             specify image size [default: 720x480]
  -t TIMESTAMP_ONOFF  timestamp on or off [default: on]
  -m MOVLENGTH        specity movie length (in sec; takes movie if specified) [default: 0]

  8035   Fri Feb 8 12:42:45 2013 nicolasSummaryGeneralPRC/arm mode matching calculations

Quote:

  The main issue is that flipping PR3 induces considerable astigmatism.

Yes, at 45degrees PR3 will only have a curvature of about 850m for the vertical mode of the beam, apparently not enough to stabilize the cavity.

  8034   Fri Feb 8 12:39:32 2013 yutaUpdateLockingPRMI work

Half-PRC at this time already have two changes from the previous half-PRC; PR2 replaced/flipped and different TM before BS.
PRMI has only one change from the previous PRMI; PR2 replaced/flipped.
This is why I wanted to try PRMI first. But we now recognized that MI alignment (including REFL and AS alignment) is tough without using the arms, I agree that we should try half-PRC first.

I don't exactly know what the situation in the Jamie's calculation, but to make the optical path length the same before and after flipping, PR2 holder have to move about n*t, where n is the substrate refractive index and t is the thickness of the mirror, towards PRM/PR3.

Quote:

The first rule of debugging is to only make one change at a time.

Also, we never talked about moving PR2 to adjust optical path length,

  8033   Fri Feb 8 11:07:07 2013 JamieSummaryGeneralPRC/arm mode matching calculations

Quote:

I would guess that either flipping PR2 or PR3 would give nearly the same effect (g = 0.9) and that flipping both makes it even more stable (smaller g). But what we really need is to see the cavity scan / HOM resonance plot to compare the cases.

The difference of 0.5% in mode-matching is not a strong motivation to make a choice, but sensitivity to accidental HOM resonance of either the carrier or f1 or f2 sidebands would be. Should also check for 2*f2 and 2*f1 resonances since our modulation depth may be as high as 0.3. Accidental 2f resonance may disturb the 3f error signals.

You would guess, and I would have guessed too, but the calculations tell the story.   Flipping both does not increase the stability.  The main issue is that flipping PR3 induces considerable astigmatism.  This is why flipping PR3 alone does not make the cavity stable.  I will do some simple calculations today that will demonstrate this effect.

But again, we should only change one thing at a time and understand that before moving on.  Given that the calculations show that flipping only PR2 should alone have a positive affect, we should do just that first, and verify that we understand what's going on, before we move on to making more changes.

I will try to make some more arbcav runs as well, for just the flipped PR2.

  8032   Fri Feb 8 11:01:18 2013 JamieUpdateLockingPRMI work

I completely agree with Koji.  We definitely should have locked the half PRC first.  We were all set up for that.  Why go through all this work to align MICH when we haven't confirmed with the half PRC that the flipping is helping us?  The first rule of debugging is to only make one change at a time.  We have measurements from the half PRC, so we could have made a direct comparison with those to see how things have changed.  If we jump the gun we're going to end up wasting more time when we have to back-track.

Also, we never talked about moving PR2 to adjust optical path length, although I can understand why we would think that should be done.  My calculations were all done assuming the free-space separation between PRM/PR2 and PR2/PR3 were unchanged.  It's possible changing the position is better, but again, it's more work and it changes multiple things at one.  I can redo my calculations for this new scenario, but we need to update our drawings with this new configuration.  Please note precisely where PR2 has moved to.

We should have just flipped PR2 and that's it.  Then we could have run the exact same measurements we had previously.  Only then, once we understood this new simple cavity, should we have done further adjustments.

  8031   Fri Feb 8 02:38:04 2013 KojiUpdateLockingPRMI work

I feel it's too hasty to use the PRMI.
I support the idea of the half-PRC test, to make an apple-to-apple comparison.

Make haste slowly.

  8030   Fri Feb 8 02:12:14 2013 JenneUpdateLockingPRMI work

[Jenne, Yuta]

Lots of work, no solid conclusions yet.  In-vac, we aligned MICH and the PRM.  Out of vac, we got beam on AS and REFL paths.  We can lock MICH, but we're not as happy with PRCL.

 

In-vac alignment: 

To get the beam centered on TT2 in yaw, Koji helped us out and moved TT1 with the sliders a little bit. Then to get the beam centered on PRM and PR2, Koji moved the TT2 sliders a little bit.

Yuta and I then moved PR2 forward a few mm, to keep the optical path length of the PRC approximately (within ~1mm, hopefully) the same as always.  After my PR2 optic swapping earlier, the pitch alignment was no longer good.  I loosened the screws holding the wire clamp to the optic holder, and tapped it back and forth until the alignment was good.  Of course, the screw-tightening / pitch-checking is a stochastic process, but eventually we got it.  A small amount of yaw adjustment by twisting the PR2 TT was also done, but not much was needed.

Beam was a little off in pitch at ITMY, so Yuta poked the top of PR3, and that one single poke was perfect, and the beam was very nicely centered on the ITMY target.  Beam was getting through BS target just fine.  We checked at ITMX, and the beam looked pretty centered, although we didn't put in a target.  We didn't do anything to BS while we were in-vac, since it was already good.

We aligned the ITMs so that their beams were retroreflecting back to the BS.  After this, we saw nice MICH fringes.

We aligned the PRM so that its beam was retroreflecting.

We checked that we were getting REFL and AS beams out of the vacuum, which we were (a small amount of adjustment was done to AS path steering mirrors).

AS table alignment:

We did a bit of tweaking of the REFL path, and lots of small stuff to the AS path. 

The AS beam was coming out of vac at a slightly different place in yaw, so we moved the first out of vac AS steering mirror so the beam hit the center, rather than ~1/3 of the way to the edge.  We then aligned the beam through the lens, to the camera, and to AS55.  Most significantly, we removed the BS that was just before AS55.  This was sending beam to a dump, but it is in place to send beam over to AS110, once we get back to real locking.  We measured ~30 microwatts of power going to the AS55 PD, while MICH fringes were fringing.

The REFL path didn't need much, although we had never been going through the center of the HWP and PBS that are used to reduce the power before going to the PDs, so we translated them a millimeter or two.

We see signal on dataviewer for all of the channels that we're interested in....AS55 I&Q, ASDC, REFL11 I&Q, REFLDC (which comes from REFL55). 

Locking:

Locking MICH was very easy, after we rotated the phase of AS55 to get all the good MICH signal in the Q phase.  Part of the criteria for this was that the AS55_Q_ERR signal should cross zero when ASDC went to 0.  This was done very coarsely, so we need to do it properly, but it was enough to get us locked.  We changed the phase from 24.5 to 90 deg.

PRCL has been more of a challenge, although we're still working on it.

On the back face of the Faraday, we see the michelson fringes, but they are not getting through the Faraday's aperture.  This implies that we have a poorly aligned michelson, in that the interference between the returning beams from the ITMs is happening at a different place than the original beam splitting.  Yuta is working on getting a better MICH right now.  EDIT, 10 minutes later....   This seems to be fixed, and the MICH fringes enter the back aperture of the FI, but there is still the PRM refl problem (next paragraph).

Also, when we get the most bright REFL beam, we see that there is some very obvious clipping in the back of the Faraday aperture, and this is matched by a clipped-looking REFL beam on the AS table.  We must understand what we have done wrong, such that when the beam is actually going through the Faraday, we see a much dimmer beam.  It's possible that there is some clipping happening at that time with the in-vac REFL path....we need to check this.  It's not a clipping problem on the AS table - I checked, and the beams are still reasonably  well centered on all of the mirrors.

We think that the MICH / REFL beam problems may be that the input pointing is close, but not perfect.  We have not confirmed today that the beam is centered on ETMY.  We should do this as part of our final alignment procedure before putting on doors.

Plans for tomorrow:

Get POP aligned, especially the camera, so we can see what our intracavity mode really looks like in the PRC.  This is probably (in part, at least) due to our having moved PR2 around, so the transmitted beams aren't in exactly the same place.

We think that it's more useful in the short term to check out the PRC, and since the clipping problem with the REFL beam is likely an imperfect input pointing, we want to use the other measured G&H mirror, and do another half-PRC test, with the test mirror in front of the BS.  This requires much less perfection in the input pointing, so it should be very quick to set up.

Confirm that PRM oplev is still aligned (turn laser back on first).

Plans for next week:

Perfect the input pointing, by checking the beam position at ETMY.  Recheck all corner alignment.

Try again locking PRMI in air.  First, confirm ITM and BS oplevs are all aligned.

  8029   Fri Feb 8 00:23:33 2013 ranaSummaryGeneralPRC/arm mode matching calculations

 

 I would guess that either flipping PR2 or PR3 would give nearly the same effect (g = 0.9) and that flipping both makes it even more stable (smaller g). But what we really need is to see the cavity scan / HOM resonance plot to compare the cases.

The difference of 0.5% in mode-matching is not a strong motivation to make a choice, but sensitivity to accidental HOM resonance of either the carrier or f1 or f2 sidebands would be. Should also check for 2*f2 and 2*f1 resonances since our modulation depth may be as high as 0.3. Accidental 2f resonance may disturb the 3f error signals.

  8028   Thu Feb 7 19:25:22 2013 yutaUpdateCDSC1ALS filters reloaded

Filters for C1ALS were all gone. So, I copied /opt/rtcds/caltech/c1/chans/C1GCV.txt and renamed it as C1ALS.txt.

I also fixed links in the medm screens; C1ALS.adl and C1ALS_COMPACT.adl.
I'm not sure what happened to C1SC{X,Y} screens.

Quote:

I decided to rename the c1gcv model to be c1als.  This is in an ongoing effort to rename all the ALS stuff as ALS, and get rid of the various GC{V,X,Y} named stuff.

(...snip...)

The above has been done.  Still todo:

  • FIX SCRIPTS!  There are almost certainly scripts that point to GC{V,X,Y} channels.  Those will have to be fixed as we come across them.
  • Fix the c1sc{x,y}/master/C1SC{X,Y}_GC{X,Y}_SLOW.adl screens.  I need to figure out a more consistent place for those screens.
  • Fix the C1ALS_COMPACT screen
  • ???

 

 

  8027   Thu Feb 7 19:24:57 2013 RijuUpdate Photodiode transimpedance

 Summary: Measurement and plot of shot-noise-intercept-current for MC REFL PD. 

Motivation:It is to measure the shot noise intercept current for MC REFL PD.

 

Result: The final plot is attached here. The plot suggests that the value of shot-noise-intercept current is 1.9mA

Discussion:

 

The plot is for the measured data of Noise voltage (V/sqrt(Hz)) vs DCcurrent(A). The fitted plot to this measured data follows the noise equation

Vnoise = gdet* sqrt[ 2e (iDC+idet)] ,  where gdet= transimpedance of the PD in RF region ~600

To get an approximate idea of the shot noise intercept current, we may follow the same procedure described in 7946 

In the present case minimum noise value is 1.46e-08 V/sqrt(Hz)

Therefore dark current(in2) ~dark noise voltage/RF transimpedance ~25pA/sqrt(Hz)

Therefore the approximate shot noise intercept current value is (25/18)^2 ~ 1.92mA, which matches well to the fitted value.

 

 

Attachment 1: reflshotnoise.pdf
reflshotnoise.pdf
  8026   Thu Feb 7 17:24:13 2013 SteveUpdateSAFETYfire extinguishers checked

The fire department weighted and pressure checked our units today. Surprisingly they found one powder filled can. We can only use HALON  gas in the lab.

 

  8025   Thu Feb 7 17:10:11 2013 KojiSummaryGeneralPRC/arm mode matching calculations

Quote:

I left out the current situation (PR2/3 with -600 RoC) and the case where only PR3 is flipped, since those are both unstable according to a la mode.

This surprises me. I am curious to know the reason why we can't make the cavity stable by flipping the PR3 as PR3 was supposed to have more lensing effect than PR2 according to my statement.

  8024   Thu Feb 7 15:46:42 2013 JenneUpdateLockingPR2 flipped

More correctly, a different G&H mirror (which we have a phase map for) was put into the PR2 TT, backwards.

Order of operations:

* Retrieve flat test G&H from BS chamber.  Put 4th dog clamp back on BS optic's base.

* Remove flat G&H from the DLC mount, put the original BS that was in that mount back.  Notes:  That BS had been stored in the G&H's clean optic box.  The DLC mount is engraved with the info for that BS, so it makes sense to put it back.  The DLC mount with BS is now back in a clean storage box.

* Remove PR2 TT from ITMX chamber.

* Remove suspension mounting block from TT frame, lay out flat, magnets up, on lint-free cloth on top of foil.

* Remove former PR2 G&H optic.

* Put what was the flat G&H test optic into the PR2 optic holder, with AR surface at the front.

* Put PR2 suspension block back onto TT frame.

* Put PR2 assembly back in the chamber, solidly against the placement reference blocks that Evan put in last Thursday.

* Close up, clean up, put labels on all the boxes so we know what optic is where.

 

Why the switcho-changeo?  We have a phase map for the G&H that is the new PR2, and a measured RoC of -706m, surface rms of 8.7nm.  Now, we can measure the former PR2 and see how it compares to our estimate of the RoC from the cavity measurements we've taken recently.

  8023   Thu Feb 7 14:10:25 2013 ManasaUpdateOpticsLaserOptik - AR Reflectivity - Bad data

Reflectivity of AR surface of LaserOptik (SN6)

 SN6_R.png

The first step measurements of R for AR surface. I am not convinced with the data....because the power meter is a lame detector for this measurement.

I'm repeating the measurements again with PDs. But below is the log R plot for AR surface.

R percentage

6000ppm @ 42 deg
3560ppm @ 44 deg
7880ppm @ 46 deg
4690ppm @ 48 deg

 

SN6_R.png

  8022   Thu Feb 7 12:56:18 2013 JamieSummaryGeneralPRC/arm mode matching calculations

NOTE: There was a small bug in my initial calculation.  The plots and numbers have been updated with the fixed values.  The conclusion remains the same.

Using Nic's a la mode mode matching program, I've calculated the PRC mode and g-parameter for various PR2/3 scenarios.  I then looked at the overlap of the resultant PRC eigenmodes with the ARM eigenmode.  Here are the results:

NOTE: each optical element below (PR2, ITM, etc.) is represented by a compound M matrix.  The z axis in these plots is actually just the free space propagation between the elements, not the overall optical path length.

ARM

This is the ARM mode I used for all comparisons:

 flat_ARM_t.pdfflat_ARM_s.pdf

  tangential sagittal
gouy shift, one-way 55.63 55.63
g (from gouy) 0.303 0.303
g (product of individual mirror g) 0.303 0.303

PRC, nominal design (flat PR2/3)

This is the nominal "as designed" PRC, with flat PR2/3 folding mirrors.

flat_PRC_t.pdfflat_PRC_s.pdf

  tangential sagittal
gouy shift, one-way 14.05 14.05
g (from gouy) 0.941 0.941
g (product of individual mirror g) 0.942 0.942

 ARM mode matching: 0.9998

PRC, both PR2/3 flipped

This assumes both PR2 and PR3 have a RoC of -600 when not flipped, and includes the affect of propagation through the substrates.

 flipped_PRC_t.pdfflipped_PRC_s.pdf

  tangential sagittal
gouy shift, one-way 19.76 18.45
g (from gouy) 0.886 0.900
g (product of individual mirror g) 0.888 0.902

ARM mode matching: 0.9806

PRC, only PR2 flipped

In this case we only flip PR2 and leave PR3 with it's bad -600 RoC:

flipped_pr2_PRC_t.pdfflipped_pr2_PRC_s.pdf

  tangential sagittal
gouy shift, one-way 18.37 18.31
g (from gouy) 0.901 0.901
g (product of individual mirror g) 0.903 0.903

ARM mode matching: 0.9859

Discussion

I left out the current situation (PR2/3 with -600 RoC) and the case where only PR3 is flipped, since those are both unstable according to a la mode.

I guess the main take away is that we get slightly better PRC stability and mode matching to the arms by only flipping PR2.

  8021   Thu Feb 7 10:35:35 2013 yutaUpdateGeneralStore optics in respective cabinets

I'm not the one who opened the ITMX table yesterday, but thanks for reminding me.
I put POP DC oscilloscope and its cables back.

Also, I relocked PMC and MC. It was unlocked since last night.

  8020   Thu Feb 7 09:03:54 2013 ManasaUpdateGeneralStore optics in respective cabinets

@Yuta

The ITMX table has been left open since yesterday. I am disconnecting your oscilloscope and closing the table.

To whomsoever it may concern...

I found about half a dozen new cvi optics (beam splitters, waveplates and lenses) lying around on the SP table.

Please store optics back in their respective cabinets if you are not using them immediately. Somebody might be looking around to use them. 

 

 

 

 

 

 

 

 

  8019   Wed Feb 6 22:39:23 2013 JamieUpdateGeneralPRC/arm mode matching with flipped PR2/PR3: coming soon

I intended to post a long analysis of the PRC/arm mode matching for the various TT situations using Nic's a la mode mode matching program, but I seem to have encountered what I think might be a bug.  I'll talk to Nic about it first thing in the AM.  Once the issue is resolved I should be able to post the full analysis fairly quickly.  Sorry about the delay.

  8018   Wed Feb 6 20:19:52 2013 ManasaUpdateOpticsG&H and LaserOptik mirrors

[Koji, Manasa]

We measured the wedge angle of the G&H and LaserOptik mirrors at the OMC lab using an autocollimator and rotation stage.

The wedge angles:

G&H : 18 arc seconds (rough measurement)

LaserOptik : 1.887 deg

  8017   Wed Feb 6 20:03:50 2013 ManasaUpdateLockingPRC cavity gains

Quote:

  Getting closer, but need to use the real measured AR reflectivity values, not the 1500 ppm guess. These should be measured at the correct angles and pol, using an NPRO.

 I'm still on that!

  8016   Wed Feb 6 20:00:06 2013 ManasaUpdateElectronicsBNC cables piled up at every corner

 [Yuta, Steve, Manasa]

There are cables piled up around the access connector area which have been victims of stampedes all the time. I have heard these cables were somehow Den's responsibility. 

Now that he is not around here:

I found piled up bnc's open at one end and with no labels lying on the floor near the access connector and PSL area. Yuta, Steve and I tried to trace them and found them connected to data channels. We could not totally get rid of the pile even after almost an hour of struggle, but we tied them together and put them away on the other side of the arm where we rarely walk.

There are more piles around the access connector...we should have a next cleanup session and get rid of these orphaned cables or atleast move them to where they will not be walked on.

109393779.png

 

 

 

  8015   Wed Feb 6 19:59:35 2013 ranaUpdateLockingPRC cavity gains

  Getting closer, but need to use the real measured AR reflectivity values, not the 1500 ppm guess. These should be measured at the correct angles and pol, using an NPRO.

  8014   Wed Feb 6 18:39:08 2013 JenneUpdateLockingPRC cavity gains

[Yuta, Jenne]

We have both calculated, and agree on the numbers for, the PRC gain for carrier and sideband.

We are using the measured arm cavity (power) loss of 150ppm....see elog 5359.

We get a PRC gain for the CARRIER (non-flipped folding) of 21, and PRC gain (flipped folding) of 20This is a 4.7% loss of carrier buildup.

We get a PRC gain for the SIDEBANDS (non-flipped folding) of 69, and PRC gain (flipped folding) of 62This is an 8.8% loss of sideband buildup.

The only difference between the "flipped" and "non-flipped" cases are the L_PR# values - for "non-flipped", I assume no loss of PR2 or PR3, but for the "flipped" case, I assume 1500ppm, as in Rana's email.  Also, all of these cases assume perfect mode matching.  We should see what the effect of poor mode matching is, once Jamie finishes up his calculation.

Why, one might ask, are we getting cavity buildup of ~20, when Kiwamu always quoted ~40?  Good question!  The answer seems, as far as Yuta and I can tell, to be that Kiwamu was always using the reflectivity of the ITM, not the reflectivity of the arm cavity.  The other alternative that makes the math work out is that he's assuming a loss of 25ppm, which we have never measured our arms to be so good.

 

For those interested in making sure we haven't done anything dumb:


ppm = 1e-6;

% ||      |      |        ||            ||
% PRM    PR2    PR3      ITM           ETM

T_PRM = 0.05637;
t_PRM = sqrt(T_PRM);
L_PRM = 0 *ppm;
R_PRM = 1 - T_PRM - L_PRM;
r_PRM = sqrt(R_PRM);

T_PR2 = 20 *ppm;
t_PR2 = sqrt(T_PR2);
L_PR2 = 1500 *ppm;
R_PR2 = 1 - T_PR2 - L_PR2;
r_PR2 = sqrt(R_PR2);

T_PR3 = 47 *ppm;
t_PR3 = sqrt(T_PR3);
L_PR3 = 1500 *ppm;
R_PR3 = 1 - T_PR3 - L_PR3;
r_PR3 = sqrt(R_PR3);

T_ITM = 0.01384;
t_ITM = sqrt(T_ITM);
L_ITM = 0;%100 *ppm;
R_ITM = 1 - T_ITM - L_ITM;
r_ITM = sqrt(R_ITM);

T_ETM = 15 *ppm;
t_ETM = sqrt(T_ETM);
L_ETM = 0 *ppm;
R_ETM = 1 - T_ETM - L_ETM;
r_ETM = sqrt(R_ETM);

rtl = 150*ppm;  % measured POWER round trip loss of arm cavities.
rtl = rtl/2;     %    because we need the sqrt of the exp() for ampl loss....see Siegman pg414.

eIkx_r = exp(-1i*2*pi);
r_cav_res = -r_ITM + (t_ITM^2 * r_ETM * eIkx_r * exp(-rtl)) / (1 - r_ITM*r_ETM * eIkx_r * exp(-rtl) );

eIkx_ar = exp(-1i*pi);
r_cav_antires = -r_ITM + (t_ITM^2 * r_ETM * eIkx_ar * exp(-rtl)) / (1 - r_ITM*r_ETM * eIkx_ar * exp(-rtl) );


%% PRC buildup gain

g_antires = t_PRM*eIkx_ar / (1-r_PRM*r_PR2*r_PR3*r_cav_antires*eIkx_ar);
G_ar = g_antires^2;
G_ar = abs(G_ar)  % Just to get rid of the imag part that matlab is keeping around.

g_res = t_PRM*eIkx_r / (1-r_PRM*r_PR2*r_PR3*r_cav_res*eIkx_r);
G_r = g_res^2;
G_r = abs(G_r)

  8013   Wed Feb 6 15:39:19 2013 SteveUpdateElectronicsDC power supplies in cabinets

 East arm cabinet E9 and E10

 

Attachment 1: IMG_0066_1.JPG
IMG_0066_1.JPG
  8012   Wed Feb 6 15:20:55 2013 yutaSummaryGeneralFWHM was wrong

I have to blame Jamie for putting extra 2 randomly.
Measured PRM-PR2 cavity finesse was actually 108 +/- 3 (even if you use digital system to get data).

Lorentzian fit:
  Lorentzian function is;

f(x;x0,gamma,A) = A * gamma**2/((x-x0)**2+gamma**2)

  where x0 is the location of the peak, gamma is HWHM, and A is the peak height.
  Lorentzian fitting function in my original code (/users/yuta/scripts/modescanresults/analyzemodescan.py) was

fitFunc = lambda p,x,m: (m-p[2])*p[0]**4/(4*(x-p[1])**2+p[0]**4)+p[2]

  In this function, p[0] is sqrt(FWHM), not sqrt(HWHM). I doubled gamma to make it FWHM and squared it because they should be positive.
  During Jamie's modification of my code, he doubled p[0]**2 to get FWHM, which is wrong (/users/jrollins/modescan/modescan.py).

  I should have commented that p[0] is sqrt(FWHM).

Redoing the analysis:
  1. I pulled 2 out, and modified Jamie's modescan.py so that you can name each peak with peakdistinguish=True option. I also modified fitpeak function so that it throws away "peaks" which don't look like a peak.

  2. If you run /users/yuta/PRCmodescan/run.py and name each peak, you will get peaks.csv which includes peak position, FWHM, and the type of the peak;

0.065017,0.001458,l
0.070446,0.001463,3
0.075940,0.001509,2
0.081552,0.001526,1
0.087273,0.001565,0
0.112027,0.001911,u
0.278660,0.002211,u
0.306486,0.001658,0
0.312480,0.001576,1
0.313626,2.507910,
0.318486,0.001626,2
0.319730,2.633097,
0.324801,0.001739,3
0.331848,0.001922,l
0.527509,0.001603,l
0.533231,0.001445,3
0.538648,0.001488,2
0.544081,0.001455,1
0.549517,0.001498,0
0.551725,2.422759,
0.570972,0.001346,u


  3. /users/yuta/PRCmodescan/calcmodescanresults.py reads peaks.csv and tells you the results;

Time between TEM00 and sideband  0.0239435  pm  0.00115999887452  sec
Calibration factor is  462.167602898  pm  22.3907907867  MHz/sec
FSR is  78.4797010471  MHz
FWHM is  0.729828720682  pm  0.0174145743828  MHz
TMS is  2.64718671684  pm  0.0538858477824  MHz
Finesse is  107.53166986  pm  2.5658325169
Cavity g-factor is  0.994390582331  pm  0.000228155661075
Cavity g-factor is  0.988812630228  pm  0.000453751681357   (Edited by YM; see elog #8056)
RoC of PR2 is  -187.384503001  pm  4.26100999578  m (assuming PRM RoC= 122.1  m)
RoC of PRM is  217.915890722  pm  5.65451518991  m (assuming PR2 RoC= -600  m)

  8011   Wed Feb 6 15:11:21 2013 JenneUpdateElectronics"Temporary" power supply situation

[Jenne, Yuta, Rana, Steve, Manasa]

We have taken stock of the lab "temporary" power supply situation, and things look much better.

This morning, I removed 2 unused power supplies and a function generator from the PSL table - these had been used for MC ringdown things.

I also removed the non-connected cables that had been used for the RAMMON setup, and the EOM temperature controller circuit.

This afternoon, Yuta removed the 2 HV power supplies that were used to keep PZT2 working near the end of its life.  Since we now have the active TTs, these have been turned off for a while, and just needed to be removed.

Manasa removed the power supply under the POP/POX table that was powering the amplifier for POP22.  If we are going to continue using a Thorlabs PD for POP22, then we need to make a twisted pair of wires (~20 feet) to get power from 1X1.  If we are going to (finally) install a gold RF PD, then that will not be necessary.

I removed the power supply sitting near the bottom of the LSC rack, for another amplifier for POP22 (with minicircuits filters attached).  Again, if we get a gold RF PD, we can remove the filters and amplifier.  If not, we can use the existing twisted pair of wires, and plug them into the rack rather than a power supply. 

The power supply under the NE corner of the PSL table was no longer in use.  It was most recently used for amplifiers for the green beat PDs, as Yuta mentioned in elog 6862, those were moved over to 1X2.  In elog 8008 I mentioned that Yuta and I moved those amplifiers over to rack power.

The HV supply, the function generator and the OSA controllers that were on top of the short OMC rack next to the AS table have all been removed.  We need to come up with a better place for the OSAs, since we need to re-install them.  The power supply and the function generator (which was used just for a voltage offset) were formerly used for the output steering PZTs, but lately we have just been using those mirrors as fixed mirrors, since we don't need to steer into the OMC.  Some day, we will replace those mirrors with the output steering active tip tilts, and re-commission the OMC....someday.

The power supply for the amplifier set (that goes with the set of minicircuits filters) for the RAMMON PD (which took light from the IPPOS path) has been removed.  If we determine that we need RAMMON back, we will have to make a twisted pair to power those amplifiers.

SUMMARY:

* If we don't install a gold RF PD for POP22, we need a 20ft twisted pair for +15/GND.

* Also, if we don't install a gold RF PD for POP22, we need to plug the amplifier at the LSC rack into the rack power (twisted pair already exists).

* If we need RAMMON back, we will need a twisted pair to power those amplifiers.

* All other power supplies have been removed, and put away.  We currently have 0 "temporary" power supplies in use in the lab!

  8010   Wed Feb 6 15:10:22 2013 SteveUpdatePEMnew safety signs on exterier doors

Quote:

Quote:

The wood exteier walls, gutters and doors were painted at CES-Annex building #69

 The east and south end of the 40m emergency exit doors are sealed- tapped off temporarily.  They will be painted on the out side only.  This job will be done by tomorrow noon

 Do not open chamber if you smell the paint !

 The east and south end of the 40m emergency exit doors received new safety signs.

Attachment 1: IMG_0068.JPG
IMG_0068.JPG
  8009   Wed Feb 6 15:05:18 2013 SteveUpdateSAFETYcameras must be anchord

Cameras must be immediately anchord to avoid a possible collusion with the view port !

Attachment 1: IMG_0070_1.JPG
IMG_0070_1.JPG
Attachment 2: IMG_0069_1.JPG
IMG_0069_1.JPG
  8008   Wed Feb 6 14:51:25 2013 JenneUpdateElectronics1 power supply replaced!

Quote:

Currently, DC power for amplifiers ZHL-1000LN+ is supplied by Aligent E3620A.
I tried to use power supply from the side of 1X1 rack, but fuse plug(Phoenix Contact ST-SI-UK-4) showed red LED, so I couldn't use it.

 Yuta, Jenne

We fixed things so that we are now using regular fused rack power for these amplifiers.  The fuse no longer had a red LED, but it measured open when we checked the resistance.  Although, somehow (magic?) 13.73V were getting to the other side of the fuse. 

Anyhow, replacing the fuse with a new one fixed the problem right up.

  8007   Wed Feb 6 11:59:12 2013 JenneUpdateLockingPRC cavity gains

EDIT:  These numbers are for a perfect, non-lossy arm cavity.  So, a half real, half imaginary world.

Carrier uses arm cavity reflectivity for perfectly resonant case.

PRC carrier gain, flipped PR2, PR3 = 61

PRC carrier gain, regular PR2, PR3 = 68  (same value, within errors, for no folding at all).

Carrier gain loss = (68-61)/68 = 10%

 

SB uses arm cavity reflectivity for perfectly anti-resonant case.

PRC SB gain, flipped PR2, PR3 = 21

PRC SB gain, regular PR2, PR3 = 22 (same value, within errors, for no folding at all). <--- yes, this this "regular PR2, PR3 = 22..."

SB % gain loss = (22-21)/22 = 4.5%

 

I claim that we will be fine, recycling gain-wise, if we flip the folding mirrors.  If we do as Yuta suggests and flip only one folding mirror, we'll fall somewhere in the middle.

  8006   Tue Feb 5 19:32:47 2013 yutaSummaryGeneralPR2/PR3 flipping and PRC stability

We are considering of flipping PR2 and/or PR3 to make PRMI stable because PR2/PR3 seems to be convex.
I calculated dependency of the PRC stability on the PR2/PR3 curvature when PR2/PR3 flipped and not flipped.
Flipping looks OK, from the stability point of view.

Assumption:
 PRM-PR2 distance = 1.91 m
 PR2-PR3 distance = 2.33 m
 PR3-ITM distance = 2.54 m
 PRM RoC = +122.1 m
 ITM RoC = Inf

 theta_inc PRM = 0 deg
 theta_inc PR2 = 1.5 deg
 theta_inc PR3 = 41 deg 
          (all numbers from elog #7989)

 Here, RoC means RoC measured from HR side. RoC measured from AR side will be -n_sub*RoC, assuming flat AR surface.
 I also assumed mirror thickness to be negligible.

Method:
  1. I used Zach's arbcav and modified it so that it only tells you your cavity is stable or not.
   (It lives in /users/yuta/scripts/mode_density_PRC/stableornot.m)

  2. Swept PR2/PR3 RoC (1/RoC from -0.005 to 0.005 1/m) to see the stability condition.

Results:
  1. Stability condition of the PRMI when PR2 and PR3 is not flipped is depicted in the graph below. Black region is the unstable region. We all know that current PRMI is unstable, so we are in the black region.
PRMI_PR2HR_PR3HR.png

  2. Stability conditions of PRMI with one of the PR2/PR3 flipped are depicted in the graphs below. If we flip one of them, PRMI will likely to be stable, but if the flipped one is close to flat and the RoC of the other one is  >~ -250 m (more convex than -250 m), PRMI will remain unstable.
PRMI_PR2AR_PR3HR.pngPRMI_PR2HR_PR3AR.png


  3. Stability condition of PRMI with both PR2 and PR3 flipped is depicted in the graph below. If we flip both, PRMI will be stable.
PRMI_PR2AR_PR3AR.png


Discussion:
  1. Flipping one of PR2/PR3 seems OK, but I cannot guarantee. TMS measurement insists RoC of PR2 to be ~ -190 m, if we believe PRM RoC = +122.1 m (elog #7997). We need more precise measurement if we need to be sure before flipping. I prefer PR2 flipping because PR3 flipping gives us longer path in the substrate and larger astigmatism. Also, PR3 RoC is phase-map-measured to be ~ -600 m and PR2 RoC seems to be more convex than -600 m from the TMS measurement.

  2. Flipping both is good from stability point of view. We need calculation of the loss in the PRC (and mode-mismatch to the arms). Are there any requirements?

  3. If we are going to flip PR3, are there any possibilities of clipping the beam at PR3? We need to check.

  4. I need to calculate whether mirror thickness and AR surface curvature are negligible or not.

Conclusion:
  I want to flip only PR2 and lock PRMI.

By the way:
 I don't like matlab plots.

  8005   Tue Feb 5 19:16:22 2013 JamieSummaryGeneralarbcav of PRC with +600 RoC PR2/3

This is just a simple rerun of arbcav from #7995 but with the PR2/3 RoCs set to 600, instead of -600.  Overall g-factor = 0.922, and the modes are well separated:

mode_density_PRC_3.pdf mode_density_PRC_2.pdf 

This doesn't take into account the effect of traveling through the substrates (still working on it).  It assumes the PR2/3 have been moved such that the cavity fold lengths remain the same.

This is something that we need to keep in mind: we will need to adjust the position of the PR2/3 to keep the fold lengths the same.

  8004   Tue Feb 5 15:31:03 2013 SteveUpdateGeneralclean assembly room benches cleaned up

Manasa, Jamie and Steve,

Tip-Tilts and parts moved into the most north " 40m "  cabinet  in the assembly room.

Green-black glass and related components were moved to the 40m E0 cabinet in plastic boxes.

The north flow bench has a few items that belong to us: HE/Ne laser, qpd on translation stages, an iris and one red mirror.  These were moved to the north edge of this bench.

However this leveled table is still full with other people's stuff

Attachment 1: IMG_0057.JPG
IMG_0057.JPG
Attachment 2: IMG_0061.JPG
IMG_0061.JPG
  8003   Tue Feb 5 12:08:43 2013 Max HortonUpdateSummary PagesUpdating summary pages

Getting started:  Worked on understanding the functionality of summary_page.py.  The problem with the code is that it was written in one 8000 line python script, with sparse documentation.  This makes it difficult to understand and tedious to edit, because it's hard to tell what the precise order of execution is without tracing through the code line by line.  In other words, it's difficult to get an overview of what the code generally does, without literally reading all of it.  I commented several functions / added docstrings to improve clarity and start fixing this problem.

Crontab:  I believe I may have discovered the cause of the 6PM stop on data processing.  I am told that the script that runs the summary_pages.py is called every 6 hours.  I believe that at midnight, the script is processing the next day's data (which is essentially empty) and thus not updating the data from 6PM to midnight for any of the days.

Git:  Finally, created git repository called public_html/__max_40m-summary_testing to use for testing the functionality of my changes to the code (without risking crashing the summary_pages).

  8002   Tue Feb 5 11:30:19 2013 KojiSummaryGeneralrough analysis of aligned PRM-PR2 mode scan

Makes sense. I mixed up n and n-1

Probability function: X = (x1 + x2 + ... + xn)/n, where xi = xavg +/- dx

Xavg = xavg*n/n = xavg

dXavg^2 = n*dx^2/n^2
=> dXavg = dx/sqrt(n)

Xavg +/- dXavg = xavg +/- dx/sqrt(n)

  8001   Tue Feb 5 10:18:54 2013 SteveUpdatePEMhigh particle count ALART

Quote:

Quote:

The BS camber is open only. We should close ASAP

Outside air quality is 1.7- 2.2  million particles  / cf min of 0.5 micron

 Air is still bad and the chambers are closed. Before lunch  Jamie repointed the PRM oplev. Manasa and I reset oplevs: BS and ITMX.

ETMX and ETMY are fine.

SRM and ITMY oplevs needs more work.

 The bad outside air quality is pushing up the inside counts.

The outside air is 5 million counts / cf min for 0.3 micron and 2 million counts / cf min for 0.5 micron particles

Do not open chamber over 10,000 counts / cf min of 0.5 micron

Attachment 1: badair.png
badair.png
Attachment 2: 100d_airq.png
100d_airq.png
  8000   Tue Feb 5 10:09:08 2013 yutaSummaryGeneralrough analysis of aligned PRM-PR2 mode scan

stdev of [0.764, 0.751] is 0.007, but what we need is the error of the averaged number. Statistical error of the averaged number is stdev/sqrt(n).

Quote:

0.764 and 0.751 do not give us the stdev of 0.005.

  7999   Tue Feb 5 09:08:11 2013 SteveUpdatePEM exterier doors to be painted

Quote:

The wood exteier walls, gutters and doors were painted at CES-Annex building #69

 The east and south end of the 40m emergency exit doors are sealed- tapped off temporarily.  They will be painted on the out side only.  This job will be done by tomorrow noon

 Do not open chamber if you smell the paint !

Attachment 1: IMG_0055.JPG
IMG_0055.JPG
  7998   Tue Feb 5 03:16:51 2013 KojiSummaryGeneralrough analysis of aligned PRM-PR2 mode scan

0.764 and 0.751 do not give us the stdev of 0.005.

I have never seen any Yokogawa in vicinity.

Quote:

 Measured FWHM was 0.764 MHz and 0.751 MHz. By taking the average, FWHM = 0.757 +/- 0.005 MHz.
 This gives you Finesse = 104 +/- 1, which is OK compared with the expectation.

What I need
 I need better oscilloscope so that we can take longer data (~1 sec) with higher resolution (~0.004 V/count, ~50kHz).
 TDS 3034B can take data only for 10 ksamples, one channel by one!  I prefer Yokogawa DL750 or later.

 

  7997   Tue Feb 5 02:04:44 2013 yutaSummaryGeneralrough analysis of aligned PRM-PR2 mode scan

I redid PRM-PR2 cavity scan using oscilloscope to avoid anti-aliasing effect.
Measured Finesse was 104 +/- 1.

Method:
 1. Splitted POP DC output into three and plugged two into oscilloscope TDS 3034B. Ch1 and Ch2 was set to 1 V/div and 20 mV/div respectively to take the whole signal and higer resolution one at the same time (Koji's suggestion). Sampling frequency was 50 kHz. Sweeping time through FWHM was about 0.001 sec, which is slow enough.
 2. Took mode scan data from the oscilloscope via network.

Preliminary results:
 Below is the plot of the data for one TEM00 peak.
PRMPR2scan.png

 The data was taken twice.
 Measured FWHM was 0.764 MHz and 0.751 MHz. By taking the average, FWHM = 0.757 +/- 0.005 MHz.
 This gives you Finesse = 104 +/- 1, which is OK compared with the expectation.

What I need:
 I need better oscilloscope so that we can take longer data (~1 sec) with higher resolution (~0.004 V/count, ~50kHz).
 TDS 3034B can take data only for 10 ksamples, one channel by one!  I prefer Yokogawa DL750 or later.

  7996   Mon Feb 4 22:46:03 2013 JamieSummaryGeneralarbcav for SRC with curved TTs

I ran Zach's arbcav on our SRC with curved TTs and the situation looks much worse than the PRC.

I used the following parameters

SRM: RoC = 142 m, T = 10%
ITM: RoC = 83.1e3 m, T = 1.4%
SRC length: 5.37 m

In this case, with TT RoC of -600, the combined cavity g-factor = 0.9986, and astigmatism from SR3 makes the cavity patently not stable.  You have to go up to an RoC of -710 before the cavity is just over the edge.

mode_density_SRC_3.pdfmode_density_SRC_2.pdf

 

  7995   Mon Feb 4 19:48:32 2013 JamieSummaryGeneralarbcav recalc of PRC with correct ITM transmission

I noticed that Koji used a high reflector for the ITMs for his full PRC arbcav calculation. I just redo it here with the correct ITM transmission and RoC for completeness.

In this case the finesse is 95, instead of 121.

mode_density_PRC_2.pdf

mode_density_PRC_3.pdf

  7994   Mon Feb 4 19:33:19 2013 yutaSummaryGeneralrough analysis of aligned PRM-PR2 mode scan

[Jenne, Yuta]

We redid PRM-PR2 cavity scan because last one (elog #7990) was taken with the sampling frequency of 2 KHz. We have also done TMS measurement.

Method:
 1. Align input TTs and PRM to align PRM-PR2 cavity.
 2. Sweep cavity length using C1:SUS-PRM_LSC_EXC.
 3. Get data using Jamie's getdata and fitted peaks using /users/jrollins/modescan/prc-pr2_aligned/run.py
 4. Calculated cavity parameters

Results:
 Below is the figure containing peaks used to do the calculation.

3peakdata.png

 From 11 MHz sidebands, calibration factor is 462 +/- 22 MHz/sec (supposing linear scan around peaks)
 FWHM is 1.45 +/- 0.03 MHz.
 TMS is 2.64 +/- 0.05 MHz.
 Error bars are statistical errors of the average over 3 TEM00 peaks.

 If we believe cavity length L to be 1.91 m, FSR is 78.5 MHz.
 So, Finesse will be 54 +/- 1 and cavity g-factor will be 0.9944 +/- 0.0002. 0.9889 +/- 0.0004   (Edited by YM; see elog #8056)
 If we believe RoC of PRM is exactly +122.1 m, measured g-factor insists RoC of PR2 to be -187 +/- 4.
 If we believe RoC of PR2 is exactly -600 m, measured g-factor insists RoC of PRM to be 218 +/- 6.

Discussion:
 1. Finesse is too small (expected to be ~100). This time, data was taken 16 KHz. Cut-off frequency of the digital antialiasing filter is ~ 5 kHz (see /opt/rtcds/rtscore/release/src/fe/controller.c). FWHM is about 0.003 sec, so it should not effect much according to my simulation.

 2. I don't know why FWHM measurement from the last one is similar to this one. The last one was taken 2 KHz, this means anti-aliasing filter of 600 Hz. This should double FWHM.

 3. Oscilloscope measurement may clear anti-aliasing suspicion.

  7993   Mon Feb 4 15:26:10 2013 JamieUpdateComputer Scripts / ProgramsNew "getdata" program to pull NDS channel data, including test points

I've added a new program called getdata (to scripts/general/getdata) that will conveniently pull arbitrary data from an NDS server, either DQ or online (ie. testpoints).

Start times and durations may be specified.  If past data is requested, you must of course be requesting DQ channels.  If no start time is specified, data will be pulled "online", in which case you can specify testpoints.

If an output directory is specified, the retrieved data will be stored in that directory in files named after the channels.  If an output directory is not specified, no output will be

Help usage:

 

controls@pianosa:~ 0$ /opt/rtcds/caltech/c1/scripts/general/getdata --help
usage: getdata [-h] [-s START] [-d DURATION] [-o OUTDIR] channel [channel ...]

Pull online or DQ data from an NDS server. Use NDSSERVER environment variable
to specify host:port.

positional arguments:
  channel               Acquisition channel. Multiple channels may be
                        specified acquired at once.

optional arguments:
  -h, --help            show this help message and exit
  -s START, --start START
                        GPS start time. If omitted, online data will be
                        fetched. When specified must also specify duration.
  -d DURATION, --duration DURATION
                        Length of data to acquire.
  -o OUTDIR, --outdir OUTDIR
                        Output directory. Data from each channel stored as
                        '.txt'. Any existing data files will be
                        automatically overwritten.
controls@pianosa:~ 0$ 

  7992   Mon Feb 4 15:06:56 2013 KojiSummaryGeneralHypothesis

Quote:

Q. How can we avoid this instability issue?
A.
1. Use flatter mirrors or at least concave mirrors.

2. Smaller incident angle to avoid emphasis of the RoC in the horizontal direction
3. Use weaker squishing force for mounting of the mirrors
4. Flip the PR3 mirror in the mounting ring by accepting the compromise that the AR surface is in the cavity.

 Another possibility is to use a ring heater to correct the curvature. I talked a bit with Aidan about this.

  7991   Mon Feb 4 11:10:59 2013 KojiSummaryGeneralrough analysis of aligned PRM-PR2 mode scan

The expected finesse is 100ish. How much can we beleive the measured number of 50?
From the number we need to assume PR2 has ~93% reflectivity.
This does not agree with my feeling that the cavity is overcoupled.
Another way is to reduce the reflectivity of the PRM but that is also unlikely from the data sheet.

The scan passed the peak in 4ms according to the fitting.
How do the analog and digital antialiasing filters affect this number?

  7990   Mon Feb 4 10:45:51 2013 JamieSummaryGeneralrough analysis of aligned PRM-PR2 mode scan

Here's a sort of rough analysis of the aligned PRM-PR2 cavity mode scan.

On Friday we took some mode scan measurements of the PRM-PR2 cavity by pushing PRM (C1:SUS-PRM_LSC_EXT) with a 0.01 Hz, 300 count sine wave.  We looked at the transmitted power on the POP DC PD and the error signal on REFL11_I.

Below is a detail of the scan, chosen because the actuation was in its linear region and there were three relatively ok looking transmission peaks with nice PDH response curves:

scan-labeled.pdf

The vertical green lines on the bottom plot indicate the rough averaged separation of the 11 MHz side-band resonances from the carrier, at +- 0.0275 s.  If we take this for our calibration, we get roughly 400 MHz / second.

The three peaks in top plot have an average FWHM of 0.00381 s.  Given the calibration above, the average FWHM = ~1.52 MHz.

If we assume a cavity length of 1.91 m, FSR = 78.5 MHz.

Putting this together we get a finesse = ~51.6.

Analysis of misaligned mode scans to follow.

  7989   Sun Feb 3 13:20:02 2013 KojiSummaryGeneralHypothesis

Rana mentioned the possibility that the PR2 curvature makes the impact on the mode stability. Entry 7988
Here is the extended discussion.


Hypothesis:

The small but non-negligible curvatures of the TT mirrors made the recycling cavity unstable or nearly unstable.


Conclusion:

If the RoC of the TT mirrors are -600 m (convex), the cavity would be barely stable.
If the RoC of the TT mirrors are less than -550m, the horizontal modes start to be unstable.
Assumption that all of the TT mirrors are concave should be confirmed.


Fact (I): Cavity stability

- The folded PRMI showed the mode stability issue. (L=6.78m from Jenne's entry 7973)
- The folded PRM-PR2-PR3-flat mirror cavity also showed the similar mode issue. (L=4.34m)
- The unfolded PRM-PR2 cavity demonstrated stable cavity modes. (L=1.91m)

Fact (II): Incident angle

- PRM 0deg
- PR2 1.5deg
- PR3 41deg

Fact (III): Mirror curvature

- RoC of PRM (PRMU02): +122.1m (measured, concave), or +115.6m (measured by the vendor)
- RoC of G&H mirrors: -600m ~ -700m (measured, I suppose the negative number means convex) (Jenne's entry 7851)
  [Note that there is no measurement of the phase map for the PR2 mirror itself.]
- RoC of LaserOptik mirrors: -625m ~ -750m (measured, I suppose that the measurement shows the mirrors are convex.) (Jan's entry 7627 and 7638)

Let's assume that the TT mirrors are always convex and have a single number for the curvature radius, say RTT


Cavity mode calculation with Zach's arbcav

1) The unfolded PRM-PR2 cavity:

The cavity becomes unstable when 0 > RTT > -122m  (This is obvious from the g-factor calculation)
==> The measured RoC of the TT mirrors predicts the cavity is stable. (g=0.98, Transverse Mode Spacing 3.54MHz)

2) The folded PRM-PR2-PR3-flat mirror cavity:

The cavity becomes unstable when RTT < -550 m
==> The measured RoC of the TT mirrors (RTT ~ -600m) predicts the cavity is barely stable (g=0.997, TMS ~600kHz).

- The instability occurs much faster than the unfolded case.
- The horizontal mode hits unstable condition faster than the vertical mode.
- The astigmatism mainly comes from PR3.

3) The folded PRMI:

The cavity becomes unstable when RTT < -550 m
==> The measured RoC of the TT mirrors (RTT ~ -600m) predicts the cavity is barely stable. (g=0.995, TMS ~500kHz)

- The instability occurs with almost same condition as the case 2)

The calculation result for the PRMI with RTT of -600 m. The code was also attached.


Q&A:

Q. What is the difference between unfolded and folded?
A. For the unfolded case, the PR2 reflect the beam only once in a round-trip.
For the folded case, each TT mirror reflects the beam twice. Therefore the lens power by the mirror is doubled.

Q. Why the astigmatism mainly comes from PR3?
A. As the angle of incidence is much bigger than the others (41deg).

Q. Why the horizontal mode is more unstable than the vertical mode?
A. Off-axis reflection of a spherical mirror induces astigmatism. The effective curvature of the mirror in
the horizontal direction
is R / Cos(theta) (i.e. longer), while it is R Cos(theta) (i.e. shorter). Indeed, the vertical and horizontal ROCs are factor of 2 different
for the 45deg incidence.

Q. Why the stability criteria for the case 2) and 3) similar?
A. Probably, once the effective curvature of the PRM-PR2-PR3 becomes
negative when RTT < -550 m.

Q. You said the case 2 and 3 are barely stable. If the TMS is enough distant form the carrier, do we expect no problem?
A. Not really. As the cavity get close to the instability, the mode starts to be inflated and get highly astigmatic.
For the case 2), the waist radii are 5.0mm and 3.7mm for the horzontal and vertical, respectively.
For the case 3), they are 5.6mm and 4.1mm for the horzontal and vertical, respectively.
(Note: Nominally the waist radius is 3.1mm)

Q. What do you predict for the stability of the PRM-PR2-Flat_Mirror cavity?
A. It will be stable. The cavity is stable until
RTT becomes smaller than -240 m.

Q. If the TT mirrors are concave, will the cavity stable?
A. Yes. Particularly if PR3 is concave.

Q. Rana mentioned the possibility that the mirrors are deformed by too tight mounting of the mirror in a ring.
Does it impact the stability of the cavity?

A. Possible. If the curvature is marginal and the mounting emphasizes the curvature, it may meet the unstable condition.

Q. How can we avoid this instability issue?
A.
1. Use flatter mirrors or at least concave mirrors.

2. Smaller incident angle to avoid emphasis of the RoC in the horizontal direction
3. Use weaker squishing force for mounting of the mirrors
4. Flip the PR3 mirror in the mounting ring by accepting the compromise that the AR surface is in the cavity.

Attachment 1: mode_density_PRC.pdf
mode_density_PRC.pdf mode_density_PRC.pdf
Attachment 2: mode_density_PRC.zip
  7988   Fri Feb 1 23:52:59 2013 ranaUpdateLockingPRM/PR2 cavity

 Very exciting result, if true. I suppose we should try to reconfirm this result by doing another phase map of PRM03.

Is it possible that PR2 is not flat? How would we test to see if the tip-tilt frame screw gives it a curvature? Perhaps we can check with COMSOL.

ELOG V3.1.3-